# Electron polarization induced by electroweak interaction in high-energy scattering off light nuclei

### v2

@@ Abstract
Electron polarization plays a significant role in studies on nuclear scattering. Nevertheless, the development of a comprehensive approach to such a problem remains challenging, particularly at the relativistic electron-energy scale. Herein, we present a theoretical approach to investigate the impact of electron polarization in scattering off unoriented light nuclei, based on the multipole expansion for the scattering cross section at high energies. Numerical calculations for stable 6,7Li and unstable 7Be nuclei show that the longitudinal polarization and weak interaction are not explicitly correlated when electrons scatter at 0 for all energy scales. In contrast, their correlation is strongly exhibited at other scattering angles when energy exceeds 10 GeV. A comparison between stable and unstable nuclei reveals that the stability of nuclei significantly affects the electron polarization contribution to the scattering cross section. This study opens new approaches to the nuclear structure problems, in particular the EMC effect, via deep inelastic electron-nucleus scattering using the unified electroweak theory.
@@
Technical advancements in generating polarized electron beams enable us to test the predictions from theoretical models and obtain additional data regarding their scattering on both oriented and unoriented nuclei [1, 2, 3, 4, 5, 6, 7, 8]. Experimentally, the polarization of an electron beam can either remain unchanged or be altered after scattering with a target [2, 5]. This means that the scattering cross section and asymmetry depend on both the magnitude of electron polarization and how it changes [2, 3, 6, 8]. One must either calculate the polarization density matrix through the cross section or precisely produce the polarized beams for experimental use [1, 4, 9, 10]. Scattering processes at high energies are often of particular interest since electrons have only longitudinal polarization in the direction of momentum transfer [2].
@@
It was initiated long ago that the multipole expansion for the electromagnetic cross section can be applied to examine the asymmetry resulting from electron polarization and nuclear spin orientation, and clarify their impact [1, 2]. A similar expansion for the electroweak one, which includes the contribution of weak interaction, has also been proposed [11, 12, 13, 14]. The nuclear form factors in the expressions in prior studies [1, 2, 11, 12] were either approximated or treated as parameters, while they were precisely calculated within later ones [13, 14] using the unified electroweak theory combined with the many-particle shell model and fractional parentage coefficients [15, 16, 17]. In particular, by employing the V-A structure of the neutral weak current along with appropriately chosen harmonic oscillator parameters, the latest study in Ref. [14] has successfully described the experimental data for the unpolarized electron scattering on 6,7Li and 7Be nuclei at MeV energies [18, 19], and provided predictive insights at GeV energies.
@@
To date, the study of high-energy electron polarization variability and its profound impact on electroweak scattering processes, which reveal information beyond what is inaccessible through pure electromagnetic interaction or unpolarized electron sources, has not been thoroughly investigated. Such a study also points towards a novel approach, based on the unified electroweak theory, for unveiling the origin of the EMC effect discovered in deep inelastic scattering (DIS) experiments [20, 21, 22]. The present Letter presents the first comprehensive investigation of all possible longitudinally polarized electron scatterings from unoriented light stable and unstable nuclei using the multipole expansion technique within the framework of the unified electroweak theory. We also examine the impact of electron polarization and its correlation with the weak interaction. This provides a basis for the future studies on asymmetry resulting solely from electron polarization and the weak interaction, once nuclear spin orientation is excluded.
The total electron-nucleus scattering cross section in the multipole expansion within the unified electroweak theory can be written as [13, 14]:
$\sigma_{\text{total}}^{\text{pol.}}(\xi,\xi^{\prime})=\sigma_{\text{total}}^{0}+\sigma_{\text{total}}^{\xi\xi^{\prime}},\vskip-20.0pt$ (1)
@@
where $\xi$ and $\xi^{\prime}$ are the electron polarizations before and after scattering, whose values are equal to +1 or 1-1- 1. The first term in Eq. (1), which does not involve electron polarization, is a sum of
@@
In Eqs. (2a)–(3c), $\varepsilon$ and $\varepsilon^{\prime}$ are the electron energies before and after scattering, respectively; $Q=K-K^{\prime}=(\omega,\mathbf{q})$ is the momentum transfer with $K=(\varepsilon,\mathbf{k})$ and $K^{\prime}=(\varepsilon^{\prime},\mathbf{k^{\prime}})$ being the electron momenta. $\alpha$ and $g$ are the electromagnetic and weak coupling constants, and $f$ is the nuclear recoil factor. $\lambda=-G_{F}m_{Z}^{2}Q^{2}/[2\sqrt{2}\pi(m_{Z}^{2}-Q^{2})],G_{F}=g^{2}/4\sqrt{2}m_{Z}^{2}\cos^{2}\theta_{\text{W}}$ with $\theta_{\text{W}}$ being the Weinberg angle, and $m_{Z}$ being the $Z^{0}$ boson mass. $g_{V}=-1/2+2x_{\text{W}},g_{A}=-1/2$ and $x_{\text{W}}\equiv\sin^{2}\theta_{\text{W}}$ are the weakinteraction parameters in the Weinberg-Salam model. The coefficients $A_{1},A_{2},B_{1},B_{2},C_{1}$, and $C_{2}$ are calculated in terms of the multipole form factors, whose explicit forms can be found in Ref. [13].
Obviously, the polarized cross sections [Eqs. (3a)–(3c)] depend on the polarization degrees $\xi$ and $\xi^{\prime}$. For simplicity, we assign $\xi\,(\xi^{\prime})=+1$ for the right-handed polarization and $\xi\,(\xi^{\prime})=-1$ for the left-handed one. We then consider all possible variations of electron polarization as following cases. Case 1: both incoming and outgoing electron beams are unpolarized. This case has been studied in detail in our recent study [14]. Case 2: electron polarizations before and after scattering are preserved, i.e., $\xi=\xi^{\prime}$. Case 3: an initially polarized beam becomes unpolarized, i.e., $\xi=+1/-1$ and $\xi^{\prime}=0$ (or $\xi^{\prime}$ is taken over +1 and 1-1- 1). Case 4: an electron beam reverses its initial polarization after scattering, i.e., $\xi=-\xi^{\prime}$.
@@
The total cross section (1) will be equivalent to that of unpolarized electron scattering when taking the summation over all possible electron polarization states. This indicates that an unpolarized electron beam has the spin components uniformly distributed in all directions [5]. We can also see that, if $\xi=\xi^{\prime}=+1/-1$, the electron polarization contribution via the pure electromagnetic term [Eq. (3a)] to the total cross section is identical to the unpolarized term [Eq. (2a)], namely, $\sigma_{F}^{11}=\sigma_{F}^{-1-1}=\sigma_{F}^{0}$. This proves that electron polarization is preserved under the electromagnetic interaction, as previously mentioned in Ref. [2]. In contrast, the contributions via the terms involving the weak interaction [Eqs. (3b)–(3c)] to the total cross section are unequal and different from the respective unpolarized ones [Eqs. (2b)–(2c)]. Particularly, as $\xi=+1/-1$ and $\xi^{\prime}=0$, only the last two terms (3b) and (3c) make a nonzero contribution to the total cross section. This reveals that electron polarization can be changed only when the weak interaction is present, while it is unchanged under the electromagnetic one. It is also easy to find a similar result for the inverse case, i.e., $\sigma_{F}^{0\pm 1}=\sigma_{F}^{\pm 10}=0$, while $\sigma_{FZ}^{0\pm 1}=\sigma_{FZ}^{\pm 10}\neq 0$ and $\sigma_{Z}^{0\pm 1}=\sigma_{Z}^{\pm 10}\neq 0$. So, we conclude that the weak interaction can make an electron beam with the $\xi=+1/-1$ initial polarization become an unpolarized beam after scattering, and vice versa. However, the total cross section (1) will vanish when the electron polarizations before and after scattering have opposite signs. This implies that the electroweak interaction cannot cause the spin orientation of all electrons to be reversed, but alters half of it. The specific expressions corresponding to all cases under discussion can be readily derived from Eqs. (3a)–(3c).
Now, we examine two typical polarized scattering cross sections, $\sigma_{\text{total}}^{\text{pol.}}(1,0)$ and $\sigma_{\text{total}}^{\text{pol.}}(1,1)$, and compare them to the unpolarized one $\sigma_{\text{total}}^{\text{unpol.}}$, where $\sigma_{\text{total}}^{\text{unpol.}}=2\sigma_{\text{total}}^{0}$ as aforementioned [13, 14, 17]. Here, $\sigma_{\text{total}}^{\text{pol.}}(1,0)$ reveals that the incoming right-handed polarization electron beam becomes unpolarized after scattering. Similarly, $\sigma_{\text{total}}^{\text{pol.}}(1,1)$ implies both incoming and outgoing electron beams being the right-handed polarization ones. Using multipole expansion for the scattering cross section, such calculations can be easily carried out once all nuclear form factors are well defined. For practical calculations, we investigate the elastic scattering of polarized electrons on ${}^{6,7}\text{Li}$ and ${}^{7}\text{Be}$ nuclei, whose form factors and other relevant parameters are taken from Ref. [14]. Numerical calculations are performed with various incident electron energies and scattering angles.
Fig. 1 shows the calculated $\sigma_{\text{total}}^{\text{pol.}}(1,0)/\sigma_{\text{total}}^{\text{unpol.}}$ ratio for 6Li, whereas Fig. 2 plots the computed $\sigma_{\text{total}}^{\text{pol.}}(1,1)/\sigma_{\text{total}}^{\text{unpol.}}$ one for 7Li. The $\sigma_{\text{total}}^{11}/\sigma_{\text{total}}^{0}$ ratio for 7Be is also displayed in Fig. 3. All the cross section ratios for three considered nuclei are listed in Tab. I. It can be seen that at energies not exceeding 101010 GeV, $\sigma_{\text{total}}^{\text{pol.}}(1,0)$ is just half of $\sigma_{\text{total}}^{\text{unpol.}}$, while $\sigma_{\text{total}}^{\text{pol.}}(1,1)$ is almost equal to the latter. A slight difference is observed when comparing the cross-section ratios for three nuclei (Tab. I). This can be attributed to the negligible contribution of weak interaction to the polarized terms at such an energy range [13]. When the electron energy goes above 10 MeV, the weak interaction gradually emerges and correlates with electron polarization to make polarized cross sections considerably increase, as seen in Figs. 1–3.
@@
We note that the exponential functions in the numerator and denominator cancel out each other when calculating all considered ratios, while the scattering cross section itself rapidly decreases toward zero with increasing scattering angles at high energies, as previously reported in Refs. [13, 14]. Therefore, their plots remain fully visible across all scattering angles. However, our data table only lists the ratios at narrow angles ($\theta=0^{\circ}-8^{\circ}$) since their values at larger angles are not particularly meaningful once all the scattering cross sections are extremely small. The results in Tab. I also indicate that the information obtained in each scattering process strongly depends on how electron polarization changes. Specifically, once electron polarization remains unchanged, the scattering cross section ($\sigma_{\text{total}}^{\text{pol.}}(1,1)$) is significantly larger than that obtained when the outgoing beam becomes unpolarized ($\sigma_{\text{total}}^{\text{pol.}}(1,0)$), in particular at energies above 100 GeV and scattering angles $\theta=5^{\circ}-8^{\circ}$.
@@
Since the unpolarized term $\sigma_{\text{total}}^{0}$ in the total polarized cross section is identical in all cases, we can make a direct comparison between $\sigma_{\text{total}}^{10}$ and $\sigma_{\text{total}}^{11}$, whose calculated ratios for three nuclei are also provided in Tab. I. At energies below 10 GeV, the $\sigma_{\text{total}}^{10}/\sigma_{\text{total}}^{11}$ ratios are very small, with the highest order of magnitude being only around $10^{-4}$ at $\theta=8^{\circ}$. However, at higher energy scales, these ratios increase remarkably, reaching up to several tens of percent. At the same energy and scattering angle, except for $\theta=0^{\circ}$, their magnitudes follow an ascending order of 7Li, 6Li, and 7Be nuclei.
To further clarify the impact of electron polarization in each scattering process, let us now consider the ratios $\sigma_{\text{total}}^{10}/\sigma_{\text{total}}^{0}$ and $\sigma_{\text{total}}^{11}/\sigma_{\text{total}}^{0}$ shown in Tab. I. Similar to the $\sigma_{\text{total}}^{10}/\sigma_{\text{total}}^{11}$ ratio, the contribution of $\sigma_{\text{total}}^{10}$ can be neglected at 10 GeV, since its highest order of magnitude relative to $\sigma_{\text{total}}^{0}$ at $\theta=8^{\circ}$ is only around $10^{-4}$. In contrast, $\sigma_{\text{total}}^{11}$ is nearly equal to $\sigma_{\text{total}}^{0}$ at energies not more than 10 GeV. At higher energy scales, the contribution of $\sigma_{\text{total}}^{10}$ becomes significant, whereas $\sigma_{\text{total}}^{11}$ even exceeds $\sigma_{\text{total}}^{0}$ by several times, as shown in Fig. 3 for 7Be.

### v3

# Polarized Electron Scattering from Light Nuclei at High Energies

@@ Abstract
We present a theoretical approach to investigate the scattering of polarized electrons from light nuclei using the multipole expansion for the scattering cross section within the framework of the unified electroweak theory. Scattering processes corresponding to different electron polarizations are analyzed and compared with the unpolarized electron scattering investigated earlier. Besides, the contribution of both polarized and unpolarized terms to the scattering cross section is examined. Numerical calculations for stable 6,7Li and unstable 7Be nuclei using the Weinberg-Salam model show that the longitudinal polarization and weak interaction are not explicitly correlated when electrons scatter at $\theta\simeq 0^{\circ}$ across all energy scales. A strong correlation emerges at the other scattering angles when the electron energy exceeds 10 GeV. This study provides additional information about nuclear structure and uncovers the role of electron polarization and its correlation with the weak interaction in each process, thus offering a more complete picture of electron-nucleus scattering.
@@
Electron polarization plays an important role in studying nuclear structure by means of scattering. Technical advancements in generating polarized electron beams [1] enable us to test the predictions from theoretical models and obtain additional data regarding their scattering on both unoriented and oriented nuclei [1, 2, 3, 4]. Experimentally, the polarization of an electron beam can either remain unchanged or be altered after scattering with a target [5, 3]. This means that the scattering cross section and asymmetry depend on both the magnitude of electron polarization and how it changes [5, 6, 7, 8, 9]. One must either calculate the polarization density matrix through the scattering cross section or precisely produce the polarized beams for experimental use [2, 1, 10, 11]. Scattering processes at high energies are often of particular interest since electrons have only longitudinal polarization along the direction of momentum transfer [5].
@@
It was initiated long ago that the multipole expansion for the electromagnetic cross section can be applied to examine the asymmetry resulting from electron polarization and nuclear spin orientation, and clarify their impact [2, 5]. A similar expansion for the electroweak cross section, which includes the contribution of weak interaction, has also been proposed [12, 13, 14, 15]. While the nuclear form factors in prior studies [2, 5, 12, 13] were either approximated or treated as parameters, they were precisely calculated in later works [14, 15] using the unified electroweak theory combined with the many-particle shell model and fractional parentage coefficients [16, 17, 18]. In particular, by employing the V-A structure of the neutral weak current along with appropriately chosen harmonic oscillator parameters, the latest study in Ref. [15] has successfully described the experimental data for the unpolarized electron scattering on 6,7Li and 7Be nuclei at MeV energies [19, 20], and provided predictive insights at GeV energies.
@@
To date, a comprehensive study of high-energy polarized electron-nucleus scattering with respect to the changeabilities of electron polarization in electroweak interaction processes, which could reveal information beyond what is inaccessible through pure electromagnetic interaction processes and unpolarized electron sources, has not been performed. This paper presents an investigation of all possible longitudinally polarized electron scatterings from unoriented light nuclei using the multipole expansion within the framework of the unified electroweak theory. It also provides a basis for future studies on the asymmetry resulting solely from the electron polarization and weak interaction, since nuclear spin orientation is excluded. Accordingly, the polarized scattering cross sections will be calculated and compared with the unpolarized scattering cross sections via their ratios. On the one hand, this enables us to compensate for the experimental limitations. Indeed, we can either perform the experiments for unpolarized or polarized electron-nucleus scattering and then deduce the result for the other case using the investigated ratios. On the other hand, the above ratios also help to clarify the contribution of electron polarization in different scattering processes when the weak interaction is involved. In addition, the impact of electron polarization and its correlation with the weak interaction are also examined through the ratios between the polarized and unpolarized terms in each scattering process.
The scattering process of high-energy polarized electrons on light nuclei can be described within the framework of the unified electroweak theory using the first-order Born approximation. Accordingly, the interaction between the electron and the nucleus, which includes electromagnetic and weak neutral current interactions, is mediated through the exchange of a virtual photon $\gamma$ and an intermediate boson $Z^{\circ}$. The differential scattering cross section is written as [14, 15]
$\sigma=\frac{4m_{e}^{2}\varepsilon^{\prime}}{f\varepsilon}\sum_{if}^{\text{---}}\left|M_{fi}\right|^{2},$ (1)
with $m_{e}$ being the electron mass, $\varepsilon$ and $\varepsilon^{\prime}$ being the electron energies before and after scattering, and $f$ being the nuclear recoil factor [5, 18]. The notation $\sum\limits_{if}^{\text{---}}$ denotes averaging over the initial spin states and summing over the final spin states for both the electron and nucleus. The scattering amplitude $M_{fi}$ is the sum of two parts corresponding to the electromagnetic interaction and the weak interaction of the form
$M_{fi}=-\frac{4\pi}{Q^{2}}[\alpha\overline{u}^{\prime}\gamma_{\mu}uJ_{F}^{\mu}(Q)+\lambda\overline{u}^{\prime}\gamma_{\mu}(g_{V}+g_{A}\gamma_{5})uJ_{Z}^{\mu}(Q)].$ (2)
Here $Q=K-K^{\prime}=(\omega,\textbf{q})$ is the 4-dimensional momentum transfer with $K=(\varepsilon,\textbf{k})$ and $K^{\prime}=(\varepsilon^{\prime},\textbf{k}^{\prime})$ being the electron momenta. $u=u(K,S)$ and $u^{\prime}=u^{\prime}(K^{\prime},S^{\prime})$ are the spinor state amplitudes of the electron before and after scattering, respectively. $S$ and $S^{\prime}$ are the vectors of electron polarization, which can be expressed approximately through the momenta by $S^{\mu}\approx\xi K^{\mu}/m_{e}$ and $S^{\prime\mu}\approx\xi^{\prime}K^{\prime\mu}/m_{e}$ with $\xi$ and $\xi^{\prime}$ being the electron polarizations or helicities, whose values are equal to +1 or 1-1 [5, 14]. $\alpha$ and $g$ are the electromagnetic and weak coupling constants. $\lambda=-G_{F}m_{Z}^{2}Q^{2}/[2\sqrt{2}\pi(m_{Z}^{2}-Q^{2})],G_{F}=g^{2}/4\sqrt{2}m_{Z}^{2}\cos^{2}\theta_{\text{W}}$ with $\theta_{\text{W}}$ being the Weinberg angle, and $m_{Z}$ being the $Z^{0}$ boson mass. $g_{V}=-1/2+2x_{\text{W}},g_{A}=-1/2$ and $x_{\text{W}}\equiv\sin^{2}\theta_{\text{W}}$ are the weak interaction parameters in the Weinberg-Salam model. $J_{F}^{\mu}(Q)$ and $J_{Z}^{\mu}(Q)$ are the electromagnetic and weak currents of the nucleus, $\mu=0,1,2,3$.
Substitute the expression (2) into (1) and perform the summation over electron states and then calculate the matrix traces, the scattering cross section depends on the contraction products in terms of $X_{\mu\nu}H_{F/FZ/Z}^{\mu\nu}$ and $Y_{\mu\nu}H_{F/FZ/Z}^{\mu\nu}$. More specifically, $X_{\mu\nu}=K_{\mu}K^{\prime}_{\nu}+K^{\prime}_{\mu}K_{\nu}+g_{\mu\nu}Q^{2}/2$ and $Y_{\mu\nu}=-i\varepsilon_{\mu\nu\alpha\beta}K^{\alpha}K^{\prime\beta}$ are the symmetric and asymmetric electron tensors respectively with $\varepsilon_{0123}=1$. The nuclear tensor $H_{F}^{\mu\nu}=\sum\limits_{if}^{\text{---}}J_{F}^{\mu*}J_{F}^{\nu}$ corresponds to the pure electromagnetic interaction, $H_{FZ}^{\mu\nu}=\frac{1}{2}\sum\limits_{if}^{\text{---}}(J_{F}^{\mu*}J_{Z}^{\nu}+J_{Z}^{\mu*}J_{F}^{\nu})$ corresponds to the interference of the electromagnetic and weak interactions, and $H_{Z}^{\mu\nu}=\sum\limits_{if}^{\text{---}}J_{Z}^{\mu*}J_{Z}^{\nu}$ belongs to the pure weak interaction. The summation and average are now restricted to the nuclear states. We then choose a cyclic coordinate system with the unit vectors as $\bm{\upzeta}_{0}=\textbf{e}_{z}, 𝛇$$\bm{\upzeta}_{\pm 1}=-(\pm\textbf{e}_{x}+i\textbf{e}_{y})/\sqrt{2}$, where three unit vectors in the Cartesian coordinate system are determined by $\textbf{e}_{z}=\underline{\textbf{q}}$, $\textbf{e}_{y}=\underline{\textbf{k}\times\textbf{k}^{\prime}}$, $\textbf{e}_{x}=\textbf{e}_{y}\times\textbf{e}_{z}$. Accordingly, the above contraction products can be further computed by expanding the nuclear transition currents in terms of the components with definite angular momenta, which are called the multipole form factors. They provide information about the structure of the nucleus along with scattering processes.
Recall that the electromagnetic current is conserved, while the longitudinal component of the weak neutral current is very small and can be neglected compared to the remaining components. As a result, applying the Wigner-Eckart theorem for the nuclear transition currents, we can obtain the expression of the scattering cross section in the form [14]
$\sigma_{\text{total}}^{\text{pol.}}(\xi,\xi^{\prime})=\sigma_{\text{total}}^{0}+\sigma_{\text{total}}^{\xi\xi^{\prime}}.$ (3)
@@
The first term in the expression (3), which does not involve electron polarization, is a sum of
@@
The coefficients $A_{1},A_{2},B_{1},B_{2},C_{1}$, and $C_{2}$ are determined in terms of the multipole form factors as follows
$A_{1}$ $=H\sum_{L}\left\{u_{C}(F_{L}^{C})^{2}+u_{T}\left[(F_{L}^{E})^{2}+(F_{L}^{M})^{2}\right]\right\},$ (6a)
$A_{2}$ $=0,$ (6b)
$B_{1}$ $=H\sum_{L}\left[u_{C}F_{L}^{C}V_{L}^{C}+u_{T}(F_{L}^{E}V_{L}^{E}+F_{L}^{M}V_{L}^{M})\right],$ (6c)
$B_{2}$ $=H\sum_{L}u_{T}^{\prime}\left(F_{L}^{E}A_{L}^{M}+F_{L}^{M}A_{L}^{E}\right),$ (6d)
$C_{1}$ $=H\sum_{L}\{u_{C}\left[(V_{L}^{C})^{2}+(A_{L}^{C})^{2}\right]+u_{T}\left[(V_{L}^{E})^{2}\right$.
$\left.+(V_{L}^{M})^{2}+(A_{L}^{E})^{2}+(A_{L}^{M})^{2}\right]\},$ (6e)
$C_{2}$ $=2H\sum_{L}u^{\prime}_{T}\left(V_{L}^{E}A_{L}^{M}+V_{L}^{M}A_{L}^{E}\right),$ (6f)
[…]
The expressions (3), (4a)-(4c) and (5a)-(5c) with the coefficients given by (6a)-(6f) are the complete multipole expansion for the high-energy polarized electron scattering cross section on light nuclei within the framework of the unified electroweak theory. The above result will return to that of the electromagnetic scattering case described by Weigert-Rose [2] when the weak interaction is not considered, as well as to that given by Willey [18] when the electron polarization is not taken into account. Numerical calculations for the case of unpolarized electron scattering cross sections [15] using the nuclear shell model with the harmonic oscillator parameter $\beta$ = 1.1526 fm-2 for $^{6}\text{Li}$ nucleus and with $\beta$ = 1.3055 fm-2 for $^{7}\text{Li}$ and $^{7}\text{Be}$ nuclei describe the experimental data in Refs. [19, 20] rather well at MeV energies. The results obtained also provide more information about the scattering processes at higher energies. Particularly, an application to the 7Li nucleus allows us to distinguish the contribution of the elastic scattering process at the ground state and that of the quasielastic scattering process in the excitation from the ground state to the nearest state, whereas one can only obtain the combined data of these two processes in experiments. However, polarized electron scattering processes have not yet been considered. So, we now assign $\xi\,(\xi^{\prime})=+1$ for the right-handed polarization and $\xi\,(\xi^{\prime})=-1$ for the left-handed polarization. We will consider all possible cases of electron polarization as follows: Both incoming and outgoing electron beams are unpolarized. This case has been studied in detail in our recent study [15]. Next, electron polarizations before and after scattering are preserved, i.e., $\xi=\xi^{\prime}$. Then, an initially polarized beam becomes unpolarized, i.e., $\xi=+1/-1$ and $\xi^{\prime}=0$ (or $\xi^{\prime}$ being taken over +1 and 1-1). This case will yield the same result as an initially unpolarized electron beam becomes polarized after scattering, since $\xi$ and $\xi^{\prime}$ play equivalent roles in expressions (5a)-(5c). Finally, an electron beam reverses its initial polarization after scattering, i.e., $\xi=-\xi^{\prime}$.
@@
It can be seen that the scattering cross section (3) will be equivalent to that of unpolarized electron scattering when we take the summation over all possible electron polarization states. This implies that an unpolarized electron beam has the spin components uniformly distributed in all directions [3]. In addition, with $\xi=\xi^{\prime}=+1/-1$, the contribution of pure electromagnetic interaction via the electron polarization term (5a) to the total differential cross section (3) is identical to that via the unpolarized term (4a) due to the coefficient $A_{2}=0$, a consequence of the multipole expansion for the nuclear transition currents. Moreover, $\sigma_{F}^{11}=\sigma_{F}^{-1-1}=\sigma_{F}^{0}$ also shows no difference between left-polarized electrons and right-polarized electrons. This relates to the fact that the pure electromagnetic interaction preserves electron polarization as previously mentioned in Ref. [5] and does not cause parity violation [9]. Meanwhile, the contributions of the polarized terms involving the weak interaction (5b)–(5c) to the total cross section are unequal and different from those of the respective unpolarized terms (4b)–(4c). In particular, when $\xi=0$ and $\xi^{\prime}=+1/-1$ (or $\xi=+1/-1$ and $\xi^{\prime}=0$), only the last two terms (5b) and (5c) make a nonzero contribution due to $\sigma_{FZ}^{0\pm 1}=\sigma_{FZ}^{\pm 10}\neq 0$ and $\sigma_{Z}^{0\pm 1}=\sigma_{Z}^{\pm 10}\neq 0$ while $\sigma_{F}^{0\pm 1}=\sigma_{F}^{\pm 10}=0$. It means that, for the cases under consideration, only the weak interaction makes electron polarization affect the scattering processes, while the electromagnetic interaction does not, regardless of whether or not the weak interaction is present. As a result, with an initially unpolarized electron beam ($\xi=0$), we can obtain a scattered beam that is either unpolarized ($\xi^{\prime}=0$) or left-/right-polarized ($\xi^{\prime}=+1/-1$) with different probabilities. In contrast, suppose we could create an initial electron beam that is completely left-/right-polarized ($\xi=+1/-1$), the multiole expansion of the scattering cross section shows that we could obtain either a scattered beam with the polarization unchanged or an unpolarized beam after scattering ($\xi^{\prime}=0$), which have nonzero probabilities. In general, the two possible processes ($\xi=0$, $\xi^{\prime}=+1/-1$) and ($\xi=+1/-1$, $\xi^{\prime}=0$) arise only when the weak interaction is present and it affects the scattering processes through the terms $\sigma_{FZ}^{0\pm 1},\sigma_{Z}^{0\pm 1}$, and $\sigma_{FZ}^{\pm 10},\sigma_{Z}^{\pm 10}$. Thus, the weak interaction may be a cause that makes an initially polarized electron beam become unpolarized after scattering, and vice versa. However, this needs to be verified experimentally through numerical calculations for specific nuclei, which will be presented in the next section below. And finally, the total differential cross section (3) will vanish when the electron polarizations before and after scattering have opposite signs. This implies that the electroweak interaction cannot cause the spin orientation of all electrons to be reversed. The specific expressions corresponding to all cases under discussion can be readily derived from the formulas (5a)–(5c).
The above analysis highlights common features of high-energy scattering processes of polarized electrons on light nuclei, elucidates the role of electron polarization and its correlation with the electromagnetic and weak interactions. Obviously, this information cannot be obtained through the multipole expansion of the scattering cross section using the electromagnetic theory or by considering unpolarized electron scattering processes. To get more detailed results, we need to calculate all the multipole form factors present in the expressions (4a)-(4c) and (5a)-(5c) of the scattering cross section (3). It can be seen that the dependence of the two components (involving and not involving electron polarization) on the multipole form factors is the same through the coefficients (6a)-(6f). However, the possible multipole components to be calculated, a finite number according to the selection rule, are different when considering specific state transitions or various nuclei. Since the multipole form factors are the reduced matrix elements of the respective operators, we first need to construct the explicit expressions for these operators, along with the formulas for calculating their matrix elements. While the calculation of single-particle matrix elements is quite simple, the calculation of matrix elements for multiparticle systems is much more complicated. In principle, we can convert multiparticle matrix elements into single-particle matrix elements for further calculation via the particle density matrix [21] or using the fractional parentage coefficient [14, 15, 18]. In particular, using the formulas for calculating the reduced matrix elements given in Refs. [14, 15], which are derived from the angular momentum theory combined with the fractional parentage coefficient method and nuclear shell model, all the multipole form factors in (6a)-(6f) can be computed directly without requiring any additional approximations. Calculations for the scattering of unpolarized electrons on 6Li and 7Li nuclei have shown good agreement with experimental data using the harmonic oscillator parameters aforementioned in the previous section. Furthermore, the convection currents in these cases are also retained. So, in the following, we will consider the scattering processes of polarized electrons on $^{6,7}\text{Li}$ and $^{7}\text{Be}$ nuclei, in a similar way as performed for the unpolarized electron cases, since the latter results have been experimentally verified. Moreover, we limit our considerations to only the elastic scattering processes when these nuclei are in their ground states. Accordingly, while the form factors of the 6Li nucleus are $F_{0}^{C}$, $F_{2}^{C}$, $F_{1}^{M}$, $V_{0}^{C}$, $V_{2}^{C}$, $V_{1}^{M}$, $A_{1}^{C}$ and $A_{1}^{E}$, the two 7Li and 7Be nuclei have more than form factors including $F_{3}^{M}$, $V_{3}^{M}$, $A_{3}^{C}$ and $A_{3}^{E}$. Generally, they all depend on the nucleon form factors and the harmonic oscillator parameter, with the specific expressions similar to those given in Ref. [15]. The nucleon form factors that describe experimental data well need to be parametrized in terms of dipoles [22, 23, 24].
@@
Now, we investigate two typical polarized scattering cross sections, $\sigma_{\text{total}}^{\text{pol.}}(1,0)$ and $\sigma_{\text{total}}^{\text{pol.}}(1,1)$, and compare them to the unpolarized one $\sigma_{\text{total}}^{\text{unpol.}}$, with $\sigma_{\text{total}}^{\text{unpol.}}=2\sigma_{\text{total}}^{0}$. Here, $\sigma_{\text{total}}^{\text{pol.}}(1,0)$ represents that the incoming right-handed polarization electron beam becomes unpolarized after scattering. Similarly, $\sigma_{\text{total}}^{\text{pol.}}(1,1)$ implies both incoming and outgoing electron beams being the right-handed polarization ones. Numerical calculations are performed with various incident electron energies and scattering angles. Fig. 1 shows the calculated $\sigma_{\text{total}}^{\text{pol.}}(1,0)/\sigma_{\text{total}}^{\text{unpol.}}$ ratio for 6Li, whereas Fig. 2 plots the computed $\sigma_{\text{total}}^{\text{pol.}}(1,1)/\sigma_{\text{total}}^{\text{unpol.}}$ one for 7Li. The $\sigma_{\text{total}}^{11}/\sigma_{\text{total}}^{0}$ ratio for 7Be is also displayed in Fig. 3. All the cross section ratios for the three nuclei under consideration are listed in Table I. It can be seen that at energies not exceeding $10$ GeV, $\sigma_{\text{total}}^{\text{pol.}}(1,0)$ is just approximately half of $\sigma_{\text{total}}^{\text{unpol.}}$, while $\sigma_{\text{total}}^{\text{pol.}}(1,1)$ is almost equal to the latter. Notably, the values of the above ratios are nearly independent of the scattering angle and remain unchanged over the MeV energy range. This means that from the experimental data of the unpolarized electron scattering cross section on light nuclei, e.g., data in Tables I and II in Ref. [20], we can deduce the values of the polarized electron scattering cross section at the respective energy and scattering angle. In contrast, if the polarized electron scattering cross section at an arbitrary energy and scattering angle is measured, we can also deduce the scattering cross section for the unpolarized case that might not have been available earlier. This gives us a more complete picture of both unpolarized and polarized electron scattering cross sections and compensates for experimental limitations. That the ratios under consideration are almost constant could be attributed to the negligible contributions of the weak interaction in the polarized terms (5b)-(5c) at such an energy range [14]. Meanwhile, when the electron energy passes 10 GeV, the above ratios strongly depend on the energy and scattering angle, except at 0. It can be explained by the fact that at such an energy scale, the weak interaction emerges and correlates with electron polarization to make polarized cross sections considerably increase, as seen in Figs. 1–3. This feature depicts the difference between pure electromagnetic scattering and electroweak scattering at high energies. A slight difference is observed when comparing the cross section ratios for three nuclei. For 6Li, since its multipole form factors have only isoscalar components, the given ratios are almost independent of the nuclear form factor when using the Weinberg-Salam model, with the parameter $\beta_{A}^{(0)}=0$. This feature is similar to the asymmetry (which is also determined by the scattering cross section ratio) in elastic scattering of electrons on the spin-0 nuclei or the nuclei with equal numbers of neutrons and protons [21]. However, it no longer holds when using other gauge models with $\beta_{A}^{(0)}\neq 0$. For the 7Li and 7Be nuclei, the results obtained depend on their multipole form factor in all gauge models because all the nuclear form factors have both isoscalar and isovector components.